# An EFT for anisotropic anti-ferromagnets: gapped Goldstones, pseudo-Goldstones, and phase transitions

### v1

@@ IV Matching to a short distance theory
$B$$\hat{\bm{x}}$$\hat{\bm{z}}$$\hat{\bm{x}}$$\hat{\bm{z}}$$(111)$${\color[rgb]{0.29,0.56,0.89}\definecolor[named]{pgfstrokecolor}{rgb}{0.29,0.56,0.89}(111)}$${\color[rgb]{0.29,0.56,0.89}\definecolor[named]{pgfstrokecolor}{rgb}{0.29,0.56,0.89}[11}{\color[rgb]{0.29,0.56,0.89}\definecolor[named]{pgfstrokecolor}{rgb}{0.29,0.56,0.89}\overline{2}}{\color[rgb]{0.29,0.56,0.89}\definecolor[named]{pgfstrokecolor}{rgb}{0.29,0.56,0.89}]}$$B$$[{11\overline{2}}]$$a$$a$
Figure 2: The crystal structure and spin arrangements in NiO below the critical temperature, $T_{\rm c}\simeq 523\text{ K}$, are characterized by a face-centered-cubic arrangements of $\rm{Ni}^{2+}$ ions. The $\rm{Ni}^{2+}$ ions align ferromagnetically along $[11\overline{2}]$ axis, indicated with bold dashed lines, lying within $(111)$ planes, depicted as colored triangular planes on the left. In this work, we consider the NiO sample in an external magnetic field, $\bm{B}$, that we choose oriented along the $[11\overline{2}]$ axis. In the following we will adopt a reference frame where the $[111]$ and $[11\overline{2}]$ axes, called hard and easy axes, coincide with the $x$- and $z$-axes respectively.
Let us now give explicit values of the effective coefficients appearing in the effective theory. To do that, we match some observable quantities obtained within the EFT with the same quantity as obtained from the short distance Hamiltonian. In particular, we will focus on the case of NiO, which is of particular phenomenological relevance (e.g., Hutchings and Samuelsen, 1971; Milano et al., 2004; Catinari et al., 2024).
In particular, the spin structure of NiO is dictated by that of its ${\rm Ni}^{2+}$ ions which, below its critical temperature, $T_{\rm c}\simeq 523$ K, are ordered in stacked parallel planes. In each of these planes, the spins are aligned ferromagnetically, with each plane being anti-ferromagnetic with respect to the adjacent ones (e.g., Rezende et al., 2019). In crystallographic jargon these are usually called $(111)$ planes — see Figure 2. Moreover, in each plane, the spins are aligned along the so-called $[11\bar{2}]$ direction, as again showed in Figure 2. The spin structure just described is dictated by the presence of two anisotropies in spin space. The first one is along the direction perpendicular to the planes mentioned above, the so-called $[111]$ direction, and is responsible precisely for forcing the spins to lie on these planes. This is the hard axis anisotropy. The second one, instead, is in the $[11\bar{2}]$ direction, and is responsible for the final orientation of the spins. This is the easy axis anisotropy.
@@
[…]
where $a$ is the cubic lattice parameter, see Figure 2. As observed by neutron scattering experiments, the next-to-nearest spin-spin coupling is by far the dominant one Hutchings and Samuelsen (1972, 1971), and the corresponding Heisenberg Hamiltonian is well approximated by,
@@
where the sum over $\langle\bm{i},\bm{j}\rangle$ runs only over lattice positions such that $|\bm{i}-\bm{j}|=a$, $J_{2}$ is the corresponding next-to-nearest neighbor coupling, and $J_{1}$, $J_{3}$, and so on are the remaining sub-leading couplings.
Let us now proceed with the matching. The magnon speed of propagation, $v_{\theta}$, can be found from the dispersion relations in Eq. (17) in the case of zero applied field, $B=0$, since, for sufficiently large momenta, they become degenerate and linear, $\omega_{q,\alpha}\simeq v_{\theta}q$. The effective coefficient $c_{1}$, instead, is nothing but the so-called perpendicular magnetic susceptibility, $c_{1}=\chi_{\perp}$, as one deduces from Eq. (9) (see also Chakravarty et al. (1988); Manousakis (1991)). Its expression in terms of the short distance parameters, instead, has been found in Esposito and Pavaskar (2023) by computing the rate of emission of one magnon by neutron scattering both within the EFT, as well as within the short distance theory, as done in Lovesey (1984). The result is,
@@
The two anisotropic couplings, $\lambda_{x}$ and $\lambda_{z}$, instead, can be determined by matching the values of the magnon gaps computed at zero magnetic field. In the short distance theory, these can be computed either by looking at the semi-classical solutions of the Landau-Lifshitz equations, or by actually quantizing the Hamiltonian and determining the spectrum of the corresponding magnon modes (e.g., Rezende et al., 2019). Both methods return the same results,
$\omega_{0,+}\big{|}_{B=0}={}$ $\mu\sqrt{2B_{E}\left(B_{Az}+B_{Ax}\right)}\,,$ (33a)
$\omega_{0,-}\big{|}_{B=0}={}$ $\mu\sqrt{2B_{E}B_{Az}}\,,$ (33b)
where $B_{E}$, $B_{Az}$ and $B_{Ax}$ are microscopic magnetic fields built out of the parameters of the Heisenberg Hamiltonian. Specifically, they are given by,
$B_{E}=\frac{2\mathcal{S}zJ_{2}}{\mu}\,,\quad B_{Ax}=\frac{2\mathcal{S}D_{x}}{\mu}\,,\quad B_{Az}=\frac{2\mathcal{S}D_{z}}{\mu}\,,$ (34)
where $z$ is the number of next-to-nearest neighbors coupled by $J_{2}$, which for NiO is $z=6$. With this at hand, the low energy anisotropic couplings are found to be,
$\lambda_{x}={}$ $\mu^{2}B_{E}B_{Ax}\,,$ (35a)
$\lambda_{z}={}$ $\mu^{2}B_{E}B_{Az}\,.$ (35b)
In Table 1 we summarize the numerical values of both the short and long distance parameters for NiO. In Figure 3, instead, we plot the gaps of the two lowest lying magnon modes of NiO as a function of the magnitude of the external magnetic field.
@@
$\mu/\mu_{\rm B}$ $\mathcal{S}$ $a$ [Å] $B_{E}$ [T] $B_{Ax}$ [T] $B_{Az}$ [T] $B_{\rm s.f.}$ [T] $v_{\theta}$ $c_{1}$ [MeV/Å] $\lambda_{x}$ [meV2] $\lambda_{z}$ [meV2]
2.18 1 4.17 968 0.635 0.011 4.62 $1.3\times 10^{-4}$ 0.58 9.80 0.17
Table 1: Summary of the relevant parameters for NiO, both for the short distance Hamiltonian and for the long distance EFT. All short distance values are taken from Rezende et al. (2019), except for the lattice parameter, $a$, which is taken from Hutchings and Samuelsen (1971). The propagation speeed, $v_{\theta}$, is fitted from the dispersion relations measured in Hutchings and Samuelsen (1971). Finally, the effective coefficients $c_{1}$, $\lambda_{x}$ and $\lambda_{z}$ are determined from the matching conditions in Eqs (32) and (35). In particular, in the expression for $c_{1}$ we traded the dominant coupling $J_{2}$ for the microscopic magnetic field $B_{E}$.
Figure 3: Gaps of the lowest lying magnon modes, $\omega_{0,\pm}$, as a function of the external magnetic field. Across the critical value, $B=B_{\rm s.f.}$, the lightest mode becomes gapless, while the higher one has a discontinuity.
Note also that the value of the critical spin flop magnetic field can also be determined in the short distance theory. In particular, it is given by Rezende et al. (2019),
$B_{\rm s.f.}\simeq\sqrt{2B_{E}B_{Az}}\,,$ (36)
where the approximation comes from considering the limit $B_{Ax,Az}\ll B_{E}$. Using Eqs. (35), this reproduces exactly what was found independently within the EFT, as discussed below Eq. (II.2).
@@ V The regime of validity of the EFT
Moreover, our theory also accounts for the effects of explicit symmetry breaking. This can be done consistently only as long as these effects are consistent with the regime of applicability of the EFT, and can therefore be included in the power counting. In detail, this can be quantified by requiring for the gaps induced by the anisotropies to be substantially smaller than the energy cutoff. This means to require that $\sqrt{\lambda_{x,z}}\ll J_{2}$. Using Eqs. (34) and (35), this can be rephrased in terms of microscopic magnetic fields, corresponding to requiring $B_{Ax,Az}\ll B_{E}$. This condition is indeed well satisfied by our prototypical anti-ferromagnet, NiO (see Table 1).
Finally, since our setup also features an external magnetic field, we also need to make sure that this is sufficiently small not to cause drastic changes is the short distance structure of the material. Similarly as above, we can estimate the maximum magnetic field, beyond which the theory loses validity, as the value for which the gap of one of the magnon modes becomes comparable to the energy cutoff. When the magnetic field becomes large, it dominates the expression for the largest gap in Eq. (14), which then become simply $\omega_{0,+}\simeq\mu B$. When this is comparable to the energy cutoff, the theory is not applicable anymore. In other words, we must further make sure that the applied magnetic field satisfies $B\ll J_{2}/\mu$. Note that, up to order one factors, this coincides with $B\ll B_{E}$.
@@ VI Outlook
In this work, we have built a low energy EFT for anisotropic anti-ferromagnets. We showed that the introduction of small anisotropies induces the emergence of a number of interesting phenomena, especially in the presence of a magnetic field, which can nonetheless be studied analytically within the regime of applicability of the EFT. This contributed to elucidate which of these phenomena are universal, meaning that they do not depend on the precise details of the short distance physics.
Moreover, an EFT description of this class of systems can help phrase the question in a language that can be understood from both a condensed matter and a high energy physics viewpoint. Such a description also makes it straightforward to compute magnon scattering amplitudes, which play an important role in determining the magnon lifetime in magnetic materials where traditional methods encounter difficulties Dyson (1956). This is also relevant in light of recent proposals to employ anti-ferromagnets as a target for light dark matter detection Esposito and Pavaskar (2023); Catinari et al. (2024); Marsh et al. (2019).
Various additional effects are, however, still left out of the present treatment. These are, for example, the inclusion of the long distance effects of dipolar interactions between spins. These are believed to be at the root of a richer spectrum observed in NiO, which features six additional light modes Milano et al. (2004). These modes all have gaps substantially below the energy cutoff of our EFT and, thus, it should be possible to systematically account for them. Furthermore, it is known that the magnon gaps can be varied by straining the anti-ferromagnetic sample Kim et al. (2022). This requires a treatment which includes interactions with phonons in our EFT. We leave these and other interesting direction for future work.
@@ VII Acknowledgements
We are grateful to Paolo Barone, Matteo Delladio, Bernard van Heck, Alberto Nicolis and Riccardo Penco for useful discussions. We are also particularly thankful to Tomáš Brauner for extensive conversations on the role of magnetic fields in anti-ferromagnets, and to José Lorenzana for very insightful exchanges on the role of discrete symmetries. The work of SP is partially supported by DOE grant DE-SC0015655.
Appendix A On the role of discrete symmetries in ferromangets and anti-ferromagnets
We now take a moment to discuss systematically the role played by discrete symmetries and, in particular, how to employ them in the low energy EFT to discriminate between ferromagnets and anti-ferromagnets. In doing so we will revisit some of the arguments that are typically found in the literature, rephrasing them in a way that is fully consistent with the proper low energy EFT ideology — see also Brauner (2024).
Consider again a bipartite lattice of spins, as discussed in Section II.1. Regardless of what magnetic system one is actually considering, the presence of long range order is indicated by the existence of a given order parameter, say $\bm{\mathcal{O}}$, which transforms linearly under the ${\rm SO}(3)$ group of spin rotations, $\bm{\mathcal{O}}\to\mathcal{R}\cdot\bm{\mathcal{O}}$, and acquires a non-zero expectation value on the ground state, $\langle\bm{\mathcal{O}}\rangle\neq 0$. This fact alone tells us that the original spin rotation symmetry is spontaneously broken to the subgroup that leaves $\langle\bm{\mathcal{O}}\rangle$ unchanged. At this level, there is no distinction between ferromagnets and anti-ferromagnets. In both instances, their low energy EFT will be formulated in terms of a 3-vector field, $\hat{\bm{n}}(x)$, such that $|\hat{\bm{n}}(x)|=1$ and $\langle\hat{\bm{n}}(x)\rangle\neq 0$. The difference arises when one wants to express the order parameter in terms of the degrees of freedom of the short distance theory, i.e. the spins of the material. In particular, one has,
@@ Appendix A On the role of discrete symmetries in ferromangets and anti-ferromagnets
where $\bm{\mathcal{M}}$ is the magnetization and $\bm{\mathcal{N}}$ is the Néel vector.
How about discrete symmetries? As mentioned again in Section II.1, both time reversal, $\mathcal{T}$, and the discrete rotation of $180^{\circ}$ in the plane containing the spins, $\mathcal{R}_{\pi}$, flip the direction of all the spins. But then, the order parameter also changes sign under either of these transformations,
@@
Again, this is true regardless of whether one is dealing with a ferromagnet or an anti-ferromagnet. It then follows that both these systems feature the same symmetry breaking pattern, i.e. ${\rm SO}(3)\times\mathcal{T}\to{\rm SO}(2)\times(\mathcal{T}\,\mathcal{R}_{\pi})$.
The rules of the game to build a low energy EFT for either of these systems are then exactly the same:
@@
Following this, the most general leading-order Lagrangian respecting these rules is the one reported also in Pavaskar et al. (2022),
@@
where $\theta(x)$ and $\phi(x)$ are the polar and azimuthal angles defining $\hat{\bm{n}}=(\sin\theta\cos\phi,\,\sin\theta\sin\phi,\,\cos\theta)$. As shown in Pavaskar et al. (2022), under an infinitesimal ${\rm SO}(3)$ rotation, the Lagrangian above changes by a total derivative, thus ensuring the invariance of the theory (modulo possible non-trivial topologies).
What discriminates a ferromagnet from an anti-ferromagnet then? The difference between the two systems lies in the behavior of the magnetization. First of all, for a ferromagnet, the magnetization changes sign under time reversal or, equivalently, under $\mathcal{R}_{\pi}$. For an anti-ferromagnet, instead, it does not. Moreover, the background value of the magnetization of a ferromagnet is non-zero, while it vanishes for an anti-ferromagnet.
Starting from the above low energy EFT, the magnetization is simply found as the temporal component of the Noether current associated to the original ${\rm SO}(3)$ symmetry, i.e. the spin density. From Eq. (38), one gets,
@@
Now it is clear that the only way that one has to reproduce the correct behavior of the magnetization described above must be if $c_{1}=0$ for ferromagnets, and $c_{3}=0$ for anti-ferromagnets, thus reducing the Lagrangian in Eq. (38) to the known ones for these two systems.
@@
In light of the discussion above, we stress what we believe is an important conceptual point. The conditions $c_{1}=0$ or $c_{3}=0$ do not follow from different symmetry breaking patterns distinguishing ferromagnets from anti-ferromagnets. At low energies they have the same symmetry breaking pattern. What is happening here is different. The magnetization in Eq. (39), which is a long distance observable, is an output of our EFT, not an input. By comparing the behavior of $\bm{s}$ under time reversal to that obtained from the microscopic theory of ferromagnets and anti-ferromagnets we are rather performing a matching procedure which determines the values of the effective coefficients, as per usual. Alternatively, if we would have included the single-time derivative term in the Lagrangian for anti-ferromagnets we would have arrived at the conclusion that they have a non-zero background magnetization, which is at odds with experiment.
Appendix B Overlap functions from “polology”

### v2

@@ IV Matching to a short distance theory
[…]
Figure 2: The crystal structure and spin arrangements in NiO below the Néel temperature, $T_{\rm N}\simeq 523\text{ K}$, are characterized by a face-centered-cubic arrangements of $\rm{Ni}^{2+}$ ions. The $\rm{Ni}^{2+}$ ions align ferromagnetically along the $11\overline{2}]$ axis, indicated with bold dashed lines, lying within $\{111\}$ planes, depicted as colored triangular planes on the left. In this work, we consider the NiO sample in an external magnetic field, H, that we choose oriented along the $11\overline{2}]$ axis. In the following we will adopt a reference frame where the $[111]$ and $11\overline{2}]$ axes, called hard and easy axes, coincide with the $x$- and $z$-axes, respectively.
Let us now give explicit values of the effective coefficients appearing in the effective theory. To do that, we match some observable quantities obtained within the EFT with the same quantity as obtained from the short distance Hamiltonian. In particular, we will focus on the case of NiO, which is of particular phenomenological relevance [e.g., 27, 17, 20].
In particular, the spin structure of NiO is dictated by that of its ${\rm Ni}^{2+}$ ions which, below its Néel temperature, $T_{\rm N}\simeq 523$ K, are ordered in stacked parallel planes. In each of these planes, the spins are aligned ferromagnetically, with each plane being antiferromagnetic with respect to the adjacent ones [e.g., 11]. In crystallographic jargon these are usually called $\{111\}$ planes — see Fig. 2. Moreover, in each plane, the spins are aligned along the so-called $11\bar{2}]$ direction, as again showed in Figure 2. The spin structure just described is dictated by the presence of two anisotropies in spin space. The first one is along the direction perpendicular to the planes mentioned above, the so-called $[111]$ direction, and is responsible precisely for forcing the spins to lie on these planes. This is the hard axis anisotropy. The second one, instead, is in the $11\bar{2}]$ direction, and is responsible for the final orientation of the spins. This is the easy axis anisotropy.
@@
$|\bm{i}-\bm{j}|={}$ $\frac{a}{\sqrt{2}}\quad\;\;\,\text{(nearest)}\,$, (30a)
$|\bm{i}-\bm{j}|={}$ $a\qquad\;\;\,\text{(next-to-nearest)}\,$, (30b)
$|\bm{i}-\bm{j}|={}$ $\sqrt{\frac{3}{2}}a\quad\text{(next-to-next-to-nearest)}\,$, (30c)
where $a$ is the cubic lattice parameter, see Fig. 2. As observed by neutron scattering experiments, the next-to-nearest spin-spin coupling is by far the dominant one [28, 27], and the corresponding Heisenberg Hamiltonian is well approximated by
@@
where the sum over $\langle\bm{i},\bm{j}\rangle$ runs only over lattice positions such that $|\bm{i}-\bm{j}|=a$, $J_{2}$ is the corresponding next-to-nearest-neighbor coupling, and $J_{1}$, $J_{3}$, and so on are the remaining sub-leading couplings.
Let us now proceed with the matching. The magnon speed of propagation, $v_{\theta}$, can be found from the dispersion relations in Eq. (17) in the case of zero applied field, ${\rm H}=0$, since, for sufficiently large momenta, they become degenerate and linear, $\omega_{q,\alpha}\simeq v_{\theta}q$. The effective coefficient $c_{1}$, instead, is nothing but the so-called perpendicular magnetic susceptibility, $c_{1}=\chi_{\perp}$, as one deduces from Eq. (9) (see also Refs. [35, 36]). Its expression in terms of the short distance parameters, instead, has been found in Ref. [6] by computing the rate of emission of one magnon by neutron scattering both within the EFT, as well as within the short distance theory, as done in Ref. [21]. The result is
@@
The two anisotropic couplings, $\lambda_{x}$ and $\lambda_{z}$, instead, can be determined by matching the values of the magnon gaps computed at zero magnetic field. In the short distance theory, these can be computed either by looking at the semiclassical solutions of the Landau-Lifshitz equations, or by actually quantizing the Hamiltonian and determining the spectrum of the corresponding magnon modes [e.g., 11]. Both methods return the same results:
$\omega_{0,+}\big{|}_{{\rm H}=0}={}$ $\mu\sqrt{2{\rm H}_{E}\left({\rm H}_{Az}+{\rm H}_{Ax}\right)}\,$, (33a)
$\omega_{0,-}\big{|}_{{\rm H}=0}={}$ $\mu\sqrt{2{\rm H}_{E}{\rm H}_{Az}}\,$, (33b)
where ${\rm H}_{E}$, ${\rm H}_{Az}$, and ${\rm H}_{Ax}$ are microscopic magnetic fields built out of the parameters of the Heisenberg Hamiltonian. Specifically, they are given by
${\rm H}_{E}=\frac{2\mathcal{S}zJ_{2}}{\mu}\,,\quad{\rm H}_{Ax}=\frac{2\mathcal{S}D_{x}}{\mu}\,,\quad{\rm H}_{Az}=\frac{2\mathcal{S}D_{z}}{\mu}\,$, (34)
where $z$ is the number of next-to-nearest neighbors coupled by $J_{2}$, the so-called “coordination number”, which for NiO is $z=6$. With this at hand, the low energy anisotropic couplings are found to be
$\lambda_{x}={}$ $\mu^{2}{\rm H}_{E}{\rm H}_{Ax}\,$, (35a)
$\lambda_{z}={}$ $\mu^{2}{\rm H}_{E}{\rm H}_{Az}\,.$ (35b)
In Table 1, we summarize the numerical values of both the short and long distance parameters for NiO. In Fig. 3, instead, we plot the gaps of the two lowest lying magnon modes of NiO as a function of the magnitude of the external magnetic field.
@@
$\mu/\mu_{\rm B}$ $\mathcal{S}$ $a$ [Å] ${\rm H}_{E}$ [kOe] ${\rm H}_{Ax}$ [kOe] ${\rm H}_{Az}$ [kOe] ${\rm H}_{\rm s.f.}$ [kOe] $v_{\theta}$ $c_{1}$ [MeV/Å] $\lambda_{x}$ [meV2] $\lambda_{z}$ [meV2]
2.18 1 4.17 9684 6.35 0.11 46.3 $1.3\times 10^{-4}$ 0.58 9.80 0.17
Table 1: Summary of the relevant parameters for NiO, both for the short distance Hamiltonian and for the long distance EFT. All short distance values are taken from Ref. [11], except for the lattice parameter, $a$, which is taken from Ref. [27]. The propagation speeed, $v_{\theta}$, is fitted from the dispersion relations measured in Ref. [27]. Finally, the effective coefficients $c_{1}$, $\lambda_{x}$ and $\lambda_{z}$ are determined from the matching conditions in Eqs. (32) and (35). In particular, in the expression for $c_{1}$ we traded the dominant coupling $J_{2}$ for the microscopic magnetic field ${\rm H}_{E}$. For the sake of the more high energy oriented reader, we recall that magnetic fields in a material are typically measured in oersted (Oe). In these units, the vacuum permeability is given by $\mu_{0}=10^{-4}$ T/Oe.
Figure 3: Gaps of the lowest lying magnon modes, $\omega_{0,\pm}$, as a function of the external magnetic field. Across the critical value, ${\rm H}={\rm H}_{\rm s.f.}$, the lightest mode becomes gapless, while the higher one has a discontinuity.
Note also that the value of the critical spin-flop magnetic field can also be determined in the short distance theory. In particular, it is given by [11],
${\rm H}_{\rm s.f.}\simeq\sqrt{2{\rm H}_{E}{\rm H}_{Az}}\,$, (36)
where the approximation comes from considering the limit ${\rm H}_{Ax,Az}\ll{\rm H}_{E}$. Using Eqs. (35), this reproduces exactly what was found independently within the EFT, as discussed below Eq. (II.2).
@@ V The regime of validity of the EFT
Moreover, our theory also accounts for the effects of explicit symmetry breaking. This can be done consistently only as long as these effects are consistent with the regime of applicability of the EFT, and can therefore be included in the power counting. In detail, this can be quantified by requiring for the gaps induced by the anisotropies to be substantially smaller than the energy cutoff. This means to require that $\sqrt{\lambda_{x,z}}\ll J_{2}$. Using Eqs. (34) and (35), this can be rephrased in terms of microscopic magnetic fields, corresponding to requiring ${\rm H}_{Ax,Az}\ll{\rm H}_{E}$. This condition is indeed well satisfied by our prototypical antiferromagnet, NiO (see Table 1).
Finally, since our setup also features an external magnetic field, we also need to make sure that this is sufficiently small not to cause drastic changes is the short distance structure of the material. Similarly as above, we can estimate the maximum magnetic field, beyond which the theory loses validity, as the value for which the gap of one of the magnon modes becomes comparable to the energy cutoff. When the magnetic field becomes large, it dominates the expression for the largest gap in Eq. (14), which then become simply $\omega_{0,-}\simeq\mu{\rm H}$. When this is comparable to the energy cutoff, the theory is not applicable anymore. In other words, we must further make sure that the applied magnetic field satisfies ${\rm H}\ll J_{2}/\mu$. Note that, up to order one factors, this coincides with ${\rm H}\ll{\rm H}_{E}$.
@@ VI Outlook
In this work, we have built a low energy EFT for anisotropic antiferromagnets. We showed that the introduction of small anisotropies induces the emergence of a number of interesting phenomena, especially in the presence of a magnetic field, which can nonetheless be studied analytically within the regime of applicability of the EFT. This contributed to elucidate which of these phenomena are universal, meaning that they do not depend on the precise details of the short distance physics.
Moreover, an EFT description of this class of systems can help phrase the question in a language that can be understood from both a condensed matter and a high energy physics viewpoint. Such a description also makes it straightforward to compute magnon scattering amplitudes, which play an important role in determining the magnon lifetime in magnetic materials where traditional methods encounter difficulties [37]. This is also relevant in light of recent proposals to employ antiferromagnets as a target for light dark matter detection [6, 20, 7].
Various additional effects are, however, still left out of the present treatment. These are, for example, the inclusion of the long distance effects of dipolar interactions between spins. These are believed to be at the root of a richer spectrum observed in NiO, which features six additional light modes [17]. These modes all have gaps substantially below the energy cutoff of our EFT and, thus, it should be possible to systematically account for them. Furthermore, it is known that the magnon gaps can be varied by straining the antiferromagnetic sample [38]. This requires a treatment which includes interactions with phonons in our EFT. We leave these and other interesting direction for future work.
@@ VII Acknowledgements
We are grateful to Paolo Barone, Matteo Delladio, Bernard van Heck, Alberto Nicolis, and Riccardo Penco for useful discussions. We are also particularly thankful to Tomáš Brauner for extensive conversations on the role of magnetic fields in antiferromagnets, and to José Lorenzana for very insightful exchanges on the role of discrete symmetries. The work of SP is partially supported by DOE grant DE-SC0015655.
Appendix A On the role of discrete symmetries in ferromagnets and antiferromagnets
We now take a moment to discuss systematically the role played by discrete symmetries and, in particular, how to employ them in the low energy EFT to discriminate between ferromagnets and antiferromagnets. In doing so, we will revisit some of the arguments that are typically found in the literature, rephrasing them in a way that, in our opinion, offers a viewpoint that is fully consistent with the proper low energy EFT ideology — see also Ref. [24].
Consider again a bipartite lattice of spins, as discussed in Sec. II.1. Regardless of what magnetic system one is actually considering, the presence of long range order is indicated by the existence of a given order parameter, say $\bm{\mathcal{O}}$, which transforms linearly under the ${\rm SO}(3)$ group of spin rotations, $\bm{\mathcal{O}}\to\mathcal{R}\cdot\bm{\mathcal{O}}$, and acquires a non-zero expectation value on the ground state, $\langle\bm{\mathcal{O}}\rangle\neq 0$. This fact alone tells us that the original spin rotation symmetry is spontaneously broken to the subgroup that leaves $\langle\bm{\mathcal{O}}\rangle$ unchanged. At this level, there is no distinction between ferromagnets and antiferromagnets. In both instances, their low energy EFT will be formulated in terms of a 3-vector field, $\hat{\bm{n}}(x)$, such that $|\hat{\bm{n}}(x)|=1$ and $\langle\hat{\bm{n}}(x)\rangle\neq 0$. The difference arises when one wants to express the order parameter in terms of the degrees of freedom of the short distance theory, i.e., the spins of the material. In particular, one has
@@ Appendix A On the role of discrete symmetries in ferromagnets and antiferromagnets
How about discrete symmetries? As mentioned again in Sec. II.1, both time reversal, $\mathcal{T}$, and the discrete rotation of $180^{\circ}$ in the plane containing the spins, $\mathcal{R}_{\pi}$, flip the direction of all the spins. But then, the order parameter also changes sign under either of these transformations,
$\bm{\mathcal{O}}\xrightarrow[]{\mathcal{T}}-\bm{\mathcal{O}}\,,\qquad\bm{\mathcal{O}}\xrightarrow[]{\mathcal{R}_{\pi}}-\bm{\mathcal{O}}\,,\qquad\bm{\mathcal{O}}\xrightarrow[]{\mathcal{T}\,\mathcal{R}_{\pi}}\bm{\mathcal{O}}\,.$ (37)
Again, this is true regardless of whether one is dealing with a ferromagnet or an antiferromagnet. It then follows that both these systems feature the same symmetry breaking pattern, i.e. ${\rm SO}(3)\times\mathcal{T}\to{\rm SO}(2)\times(\mathcal{T}\,\mathcal{R}_{\pi})$.
The rules of the game to build a low energy EFT for either of these systems are then exactly the same.
@@
Following this, the most general leading-order Lagrangian respecting these rules is the one reported also in [10],
@@
where $\theta(x)$ and $\phi(x)$ are the polar and azimuthal angles defining $\hat{\bm{n}}=(\sin\theta\cos\phi,\,\sin\theta\sin\phi,\,\cos\theta)$. As shown in Ref. [10], under an infinitesimal ${\rm SO}(3)$ rotation, the Lagrangian above changes by a total derivative, thus ensuring the invariance of the theory (modulo possible non-trivial topologies).
What discriminates a ferromagnet from an antiferromagnet then? The difference between the two systems lies in the behavior of the magnetization. First of all, for a ferromagnet, the magnetization changes sign under time reversal or, equivalently, under $\mathcal{R}_{\pi}$. For an antiferromagnet, instead, it does not. Moreover, the background value of the magnetization of a ferromagnet is non-zero, while it vanishes for an antiferromagnet.
Starting from the above low energy EFT, the magnetization is simply found as the temporal component of the Noether current associated to the original ${\rm SO}(3)$ symmetry, i.e., the spin density. From Eq. (38), one gets
@@
Now it is clear that the only way that one has to reproduce the correct behavior of the magnetization described above must be if $c_{1}=0$ for ferromagnets, and $c_{3}=0$ for antiferromagnets, thus reducing the Lagrangian in Eq. (38) to the known ones for these two systems.
@@
In light of the discussion above, we stress what we believe is an important conceptual point. The conditions $c_{1}=0$ or $c_{3}=0$ do not follow from different symmetry breaking patterns distinguishing ferromagnets from antiferromagnets. At low energies they have the same symmetry breaking pattern. What is happening here is different. The magnetization in Eq. (39), which is a long distance observable, is an output of our EFT, not an input. By comparing the behavior of $\bm{s}$ under time reversal to that obtained from the microscopic theory of ferromagnets and antiferromagnets, we are rather performing a matching procedure which determines the values of the effective coefficients, as per usual. Alternatively, if we would have included the single-time derivative term in the Lagrangian for antiferromagnets, we would have arrived at the conclusion that they have a non-zero background magnetization, which is at odds with experiment.
Appendix B Background and quadratic theory for a more general magnetic field
It is interesting to study what happens to the ground state of our theory, and to its quadratic Lagrangian, in the more general case of a magnetic field in the anisotropy plane, i.e., for $\text{\bf H}={\rm H}\cos\theta\,\hat{\bm{x}}+{\rm H}\sin\theta\,\hat{\bm{z}}$. For simplicity we limit ourself to the range $\theta\in[0,\pi/2]$. In this instance, the Hamiltonian obtained from Eq. (11), computed on static and homogeneous configurations, reads
[…]
where again we omit irrelevant constant terms. One can verify that the background configuration that minimizes the Hamiltonian depends on the value taken by the projection of the magnetic field along the easy axis. In particular, for ${\rm H}\sin\theta>{\rm H}_{\rm s.f.}$ the background is $\langle\hat{\bm{n}}\rangle=\hat{\bm{y}}$, while for ${\rm H}\sin\theta<{\rm H}_{\rm s.f.}$ it is $\langle\hat{\bm{n}}\rangle=\hat{\bm{z}}$.
In this more general case, the quadratic Lagrangian is always non-diagonal, in each of the two phases. The Fourier space linear equations of motion for the magnon fields are again given by $\mathcal{M}^{ab}(\omega,q)\theta^{b}(\omega,\bm{q})$, where now the kinetic matrix is given by
$\mathcal{M}^{ab}={}$ $\begin{pmatrix}v_{\theta}^{2}q^{2}-\omega^{2}+\mu^{2}{\rm H}^{2}\sin^{2}\theta-2\lambda_{z}&-\mu^{2}{\rm H}^{2}\cos\theta\sin\theta\\ -\mu^{2}{\rm H}^{2}\cos\theta\sin\theta&v_{\theta}^{2}q^{2}-\omega^{2}+\mu^{2}{\rm H}^{2}\cos^{2}\theta+2\lambda_{x}\end{pmatrix}\,,\qquad\qquad\quad\,\text{ for }{\rm H}\sin\theta>{\rm H}_{\rm s.f.}\,$, (41a)
$\mathcal{M}^{ab}={}$ $\begin{pmatrix}v_{\theta}^{2}q^{2}-\omega^{2}-\mu^{2}{\rm H}^{2}\sin^{2}\theta+2\lambda_{z}&2i\mu{\rm H}\omega\sin\theta\\ -2i\mu{\rm H}\omega\sin\theta&v_{\theta}^{2}q^{2}-\omega^{2}+\mu^{2}{\rm H}^{2}\cos 2\theta+2(\lambda_{x}+\lambda_{z})\end{pmatrix}\,,\qquad\text{ for }{\rm H}\sin\theta<{\rm H}_{\rm s.f.}\,.$ (41b)
Now, while neither of them is diagonal, the off-diagonal terms for the ${\rm H}\sin\theta>{\rm H}_{\rm s.f.}$ case do not depend on frequency. This means that the quadratic theory can be diagonalized by a simple, local field redefinition. In other words, one can simply work with a linear combination of the magnon fields, chosen to diagonalize the matrix. For ${\rm H}\sin\theta<{\rm H}_{\rm s.f.}$, instead, the off-diagonal terms are frequency dependent, and one thus need to follow the same procedure discussed in Sec. III.
The spectrum of the theory is found again by requiring that the matrices above have vanishing determinant. The final expressions can be found analytically, but they are considerably more involved than what was found in the main text. We thus omit them.
Appendix C Overlap functions from “polology”